In an effort to save money for early retirement, an environmental engineering colleague plans to deposit $2,000 per month, starting one month from now, into a fixed rate account that pays 8% per year compounded semi-annually. How much will be in the account at the end of 20 years?

At the end of 20 years, the account will be $
Answer #1

Deposits in one semiannual period = 2000 * 6 = 12000

i = 8% = 8% /2 = 4% per semi annual period

t = 20 * 2 = 40 semi annual period

F = 12000 * (F/A, 4%, 40)

= 12000 * 95.025516

= 1140306.19 ~ 1140306 (Nearest Dollar)